OpenStudy (anonymous):

HELP ME PLZ Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3

OpenStudy (anonymous):

when x=3 u have to substitute 3 in place of x so equation becomes 3-7(3)+2(3)=2+21+6=29

OpenStudy (anonymous):

sorry it is 3-21+6=24

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

no becauseu cant do 3-21+6

OpenStudy (anonymous):

3-21 is negitive

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its ok thank u so much

OpenStudy (anonymous):

ur welcome